Ten gallons of a 25% salt solution are to be obtained by mixing solutions that are 13% salt and 33% salt respectively. How many gallons of each are needed?
Answer by stanbon(75887)

Ten gallons of a 25% salt solution are to be obtained by mixing solutions that are 13% salt and 33% salt respectively. How many gallons of each are needed?
--------
Let amount of 13% solution be "x" ; Amt of salt is 0.13x gals.
Amt of 33% solution is "10-x" gallons: Amt of salt is 0.33(10-x)=3.3-0.33x gals
----------------------------
Amt of 25% solution is 10 gals ; Amt of salt is 0.25*10 = 2.5 gals
----------
EQUATION:
salt + salt = salt
0.13x +3.3-0.33x =2.5
-0.20x=-0.8
x= 4 gallons (amt of 13% solution to be mixed)
10-x= 6 gallons (amt of 33% solution to be mixed)
-----------
Cheers,
Stan H.